OpenStudy (iwanttogotostanford):

the transformation is a parabola on the x axis going up 7

OpenStudy (iwanttogotostanford):

@jim_thompson5910 @zepdrix

jimthompson5910 (jim_thompson5910):

what are some known points on \(\Large y = \log_2(x)\) ?

OpenStudy (iwanttogotostanford):

im not sure, I'm lost on this problem!:-( @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok let's try plugging in x values what is y when x = 1 ?

jimthompson5910 (jim_thompson5910):

you'll need the change of base formula \[\Large \log_b(x) = \frac{\log(x)}{\log(b)}\]

OpenStudy (iwanttogotostanford):

@jim_thompson5910 how do i do that please i really need help

jimthompson5910 (jim_thompson5910):

what kind of calc do you have?

OpenStudy (anonymous):

I think this is 2/3

OpenStudy (anonymous):

Hmm.

OpenStudy (anonymous):

Try f(x + 2) + 3.

OpenStudy (iwanttogotostanford):

how do i try it

OpenStudy (anonymous):

What do you mean?

OpenStudy (iwanttogotostanford):

@j.c.b

OpenStudy (anonymous):

Yes

OpenStudy (iwanttogotostanford):

thank you !

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Answer: f(x+2)+3